Consider two numbers. One of the numbers is 19 more than the other. The sum of their square is 185. Find the numbers

Found 2 solutions by mananth, jee8565:
Answer by mananth(16946)

let the numbers be x & y
x-y=19
x=y+19
x^2+y^2=185
substitute x in the second equation
(y+19)^2+y^2=185
2y^2+38y+361=185
2y^2+38y+176=0
2y^2+22y+16y+176=0
2y(y+11)+16(y+11)=0
(2y+16)(y+11)=0
2(y+8)(y+11)=0
y=-8 OR -11
The numbers are (11,-8) , (8,-11)
m.ananth@hotmail.ca

Answer by jee8565(1)

(19+(19-x))^2=185
the answer is 19-x